OpenStudy (anonymous):

f(x) = 5ln(6+x) find f'( x )

OpenStudy (anonymous):

\[f'(x)=\frac{5}{6+x}\]

OpenStudy (anonymous):

f'(x) = 5 ( 6 + x)' / (6 + x) = 5/ ( 6 + x)